Find the area of the region enclosed by the curves y= x3 - 2x2 and y = 2x2 - 3x


1
Expert's answer
2021-04-14T14:52:20-0400

Ans:-

The area between curves is the area between a curve f(x) and a curve g(x) on an interval [a, b]

"A=\intop^b_a \mid f(x)-g(x)\mid dx"

here "f(x)= x^3 - 2x^2" and "g(x)= 2x^2 - 3x"


So applying the formula "\intop^3_0 \mid x^3-2x^2-(2x^2-3x)\mid dx"


"\to \intop^3_0 \mid x^3-4x^2+3x\mid dx"


"[\dfrac{x^4}{4}-\dfrac{4x^3}{3}+\dfrac{3x^2}{2}]^3_0"


"\dfrac{37}{12} cm^2"